A cubic box measures 37.8 cm on an edge. What is the volume of the box in cubic centimeters? Express the answer to the correct number of significant figures.
1
Expert's answer
2020-06-02T14:09:07-0400

"Volume=37.8^3 cc=54010.152 cc"

Now, the least number of signinficant digits in the multiplication operation are 3, thus result must also contain 3 significant digits only.

Hence, Volume "=54.0*10^3cc"